OpenStudy (anonymous):

. Write two related equations for x + 7 = 19

OpenStudy (anonymous):

can some one help me

OpenStudy (dangerousjesse):

\(\large\textsf{19-7=x}\) \(\large\textsf{7+x=19}\) \(\large\textsf{19-12=7}\) \(\large\textsf{x=7+5}\) etc etc